# Topological excitations in a spin-orbit coupled spin-1 Bose-Einstein condensate under sinusoidally varying magnetic fields

### v1

@@ 3.2.1 Case-\@slowromancapi@ Sine-Sine magnetic field $B(\mathbf{r})=B_{0}(\rm{sin}x,\rm{sin}y)$
Figure 8: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the cosine-cosine magnetic field $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
@@
Figure 9: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 8. The unit length is the oscillator length $l$.
Figure 10: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the sine-cosine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
vortices or anti-vortices, their corresponding topological charge distributions are also elliptical. We show in Fig. 9 ($c$), the local amplification of the spin texture for the region of the diagonal skyrmion and anti-skyrmion.
@@
It is paramount to note that in the absence of SOC, the ground state structures for the $sine-sine$ case and the $cosine-cosine$ case can be analogously explained by taking into account the fact that the topological excitations are observed for certain saddle points in the magnetic fields. In this light, it becomes interesting to study the ground state structures for cross-fields, sine-cosine and/or cosine-sine fields.
3.2.3 Case-\@slowromancapiii@ Sine-Cosine magnetic field $B(\mathbf{r})=B_{0}(\rm{sin}x,\rm{cos}y)$
In this case, we study the ground state structures for the spin-1 BEC with the external magnetic field chosen as $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with equal magnetic strength $B_{0}=0.25$. For
@@ 3.2.3 Case-\@slowromancapiii@ Sine-Cosine magnetic field $B(\mathbf{r})=B_{0}(\rm{sin}x,\rm{cos}y)$
Figure 11: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-cosine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
this combination of magnetic fields, in the absence of SOC, the ground state structures exhibit density modulations due to the resultant interaction of the sine-cosine magnetic fields in the $x,y$ directions, respectively. We present the ground state density profiles of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components in Fig. 10 ($a1-a3$). In the absence of SOC, although the overall angular momentum is conserved, signature of topological excitations in the form of vortices and anti-vortices are observed in the $m_{F}=1$ and $m_{F}=-1$ components. These are arranged in pairs of alternating vorticity along the $y-$direction and as alternating vortex-anti-vortex triads in the $x-$direction. Due to different magnetic fields in different directions, the local high densities in the $m_{F}=0$ component (which arrise as an outcome of the spin-exchange interaction of $m_{F}=0$ component with the $m_{F}=\pm 1$ components), are anisotropic in nature. The vorticity profile (see Fig. 10 ($b1-b3$)) confirms the complementary response of the $m_{F}=1$ and $m_{F}=-1$ components to the external magnetic field. The momentum distribution is shown in Fig. 10 ($c1-c3$). For this case also, along with the central peak, extra momentum both in $k_{x},k_{y}$ directions is observed. However, as their magnitudes are not same in orthogonal directions, the density modulations have different wavelengths along different axes.
In the presence of spin-orbit coupling, the system experiences symmetry breaking, due to which an anisotropy is introduced in the spin-1 system under the influence of magnetic field. For this case, the three components undergo density modulations in both directions, however differently for different components. The density as well as vortex distributions are highly anisotropic in nature due to the combined effect of SOC-induced anisotropy and different sinusoidal fileds in different directions. We present the component-wise density profiles in Fig. 11 ($a1-a3$). For the $m_{F}=1$ component, there is a local high density surrounded by vortices in the system, while for the $m_{F}=-1$ component, there is a vortex surrounded by adjacent local densities. The $m_{F}=0$ component has a unique configuration for which a local high density is observed, which is surrounded by anti-vortices along the $x-$direction and an adjacent vortex along the $y-$direction. The corresponding vorticity profiles are given in Fig. 11 ($b1-b3$). For this case also, the coreless vortices have deformed elliptical shape as a consequence of induced anisotropy in the system. The momentum distributions as shown in Fig. 11 ($c1-c3$) show the effect of anisotropy in the density modulations in the spin-1 system. The k-space densities are elongated in opposite directions for the $m_{F}=1$ and $m_{F}=-1$ components, respectively, i.e, for the $m_{F}=1$ case, the central peak is slightly elongated along the $x-$axis while for the $m_{F}=-1$ component the central density peak is slightly elongated in the $y-$direction. This is to respect the density modulations with different wavelengths in the anisotropic system. In addition to this, the extra momenta in the $k_{x},k_{y}$ directions are not equivalent. For the $m_{F}=0$ component, the momentum distribution is symmetric however not the same about the $k_{x},k_{y}$ axis.
The corresponding topological charge distribution is shown in Fig. 12 ($a$). The presence of skyrmions and anti-skyrmions is confirmed by the spin texture of the system shown in Fig. 12 ($b$). Because of the combined effect of anisotropy and different magnetic fields along different directions, the spin texture has both skyrmions and anti-skyrmions. Fig. 12 ($c$) shows the local amplification of the spin texture showcasing the central vortex-anti-vortex triad.
Figure 12: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 11. The unit length is the oscillator length $l$.
3.2.4 Case-\@slowromancapiv@ Cosine-Sine magnetic field $B(\mathbf{r})=B_{0}(\rm{cos}x,\rm{sin}y)$
For the present case, we propose an alternate choice of cross-magnetic field in the $x-y$ plane, i.e., $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{sin}x$. Here, we attempt to highlight the striking similarities and subtle differences of this present case with the previous case (discussed in section 3.2.3). It is trivial to understand that for this case also, the ground state structures are direction-dependent
@@ 3.2.4 Case-\@slowromancapiv@ Cosine-Sine magnetic field $B(\mathbf{r})=B_{0}(\rm{cos}x,\rm{sin}y)$
Figure 13: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-cosine magnetic field $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
as a result of choosing different magnetic fields in different directions. Importantly, as the magnetic fields along the $x,y-$directions are swapped, their effect on the spin-1 ground state will therefore, be also swapped. In the absence of SOC, the density profiles, associated vorticities and momentum space densities of the three components are interchanged along the $x-y$ axes.
However, for the SO-coupled spin-1 BEC in the cosine-sine magnetic field, the ground state structures are not simply the transpose of that of the sine-cosine case. The ground state structure which is an outcome of the intricate interplay between the external magnetic field, SOC and the spinor order parameter is being presented in Fig. 14 for the SOC strength $\gamma=0.5$. The
@@
Figure 14: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 13. The unit length is the oscillator length $l$.
@@
SOC results in rotational symmetry-breaking of the system rendering the system an induced anisotropy. The density profiles (see Fig. 13 ($a1-a3$)) vouch for the experienced anisotropy in the system. The $m_{F}=1$ component has a unique configuration of vortices along with adjacent local high densities in the system. The $m_{F}=0$ component has a linear chain of vortices in the $y-$direction with an additional anti-vortex in the $x-$direction. The $m_{F}=-1$ component has a local high density surrounded by vortex structures on adjacent sides. The vorticity profile as shown in Fig. 13 ($b1-b3$) depicts the nature of this vortices or anti-vortices in the system. The coreless vortices have a slightly distorted shape due to the symmetry breaking. The momentum distributions are presented in Fig. 13 ($c1-c3$). Due to the induced anisotropy in the system, the k-space densities have visible elongations along a particular direction. In addition, there exists certain extra momenta in the $k_{x},k_{y}$ direction which is not symmetric.
@@
The topological charge distribution is shown in Fig. 14 ($a$). The skyrmions and anti-skyrmions are observed in different axes. The spin texture is shown in Fig. 14 ($b$). Aside from the central triad formed with alternating vortex-anti-vortex configuration, a few other spin textures can be observed owing to the coreless vortices in the system. Fig. 14 ($c$) shows the local amplification of the spin texture showing the central triad in the system.
4 Conclusion

### v2

@@ III.2.1 Case-\@slowromancapi@: Sine-Sine magnetic field ${B}(\mathbf{r})=B_{0}(\sin x,\sin y)$
Figure 8: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the cosine-cosine magnetic field $B_{x}=B_{0}\cos x$ and $B_{y}=B_{0}\cos y$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
III.2.2 Case-\@slowromancapii@: Cosine-Cosine magnetic field ${B}(\mathbf{r})=B_{0}(\cos x,\cos y)$
Here, we choose the magnetic fields as cosine functions in both directions, i.e., ${B_{x}}=B_{0}\cos x$ and ${B_{y}}=B_{0}\cos y$. The component-wise density profiles of the ground state due to the magnetic field independent of the SO-coupling are shown in Fig. 7 (a1-a3). Comparable to the sine-sine magnetic field case, here also a certain stable saddle-point structure is formed due to the interaction of periodic magnetic fields in both directions. Due to the presence of a saddle-point structure, there emerges a configuration wherein an equal number of vortices and anti-vortices are formed within the $m_{F}=1$ and $m_{F}=-1$ components, respectively. This arrangement ensures the preservation of angular momentum conservation within the spin-1 system. The spin-exchange interaction along with the effect of magnetic field, yields vortex brights in the $m_{F}=0$ component. These local high densities in the $m_{F}=0$ component compensate for the vortices or anti-vortices in the adjacent $m_{F}=\pm 1$ components. The vorticity profiles associated with the component densities are presented in Fig. 7 (b1-b3), which clearly explains the angular momentum conservation in the system. The topological excitations arise in the system in the form of exotic vortex-anti-vortex configuration at the saddle-points of the external magnetic field. The density modulations due to such exotic structures can be explained by the k-space density profiles (see Fig. 7 (c1-c3)). For this case also, apart from the central peak, there exists extra density peaks in the momentum space both in $k_{x},k_{y}$ conforming the density modulations with wavlength inversely proportional to non-zero momentum peaks in both directions. The topological charge is zero for the non SOC case to ensure the angular momentum conservation.
However, the coopetition between the cosine-cosine magnetic field and the SOC results in the re-distribution of particles in the system. The density profiles of the SO-coupled spin-1 BEC are shown in Fig. 8 (a1-a3). Trivially, the spin-mixing interaction ensures that the density profiles of $m_{F}=1$ and $m_{F}=-1$ components complement each other, with the presence of vortices in one at the loci of high density for the other component. Moreover, the vortices in one are compensated by the anti-vortices in the other. The $m_{F}=0$ component experiences topological excitations in the form of vortices arranged at the diagonal saddle-points along with local high densities at the opposite diagonal. Also, the criss-cross vortex and vortex-brights in the $m_{F}=0$ component complements the local high densities and vortices in the $m_{F}=\pm 1$ components. The vorticity profiles as shown in Fig. 8 (b1-b3) give a clear picture about the topological excitations in the system. Remarkably, the magnetic field influences the vortices in the $m_{F}=0$ component to have opposite vorticity in the ground state. In Fig. 8 (c1-c3), we present the k-space densities of the three components of the spin-1 BEC. Similar to the previously discussed cases, here also, apart from the central peak, extra density peaks along the $k_{x},k_{y}$ directions are observed, which corresponds to the density modulations. Interestingly, for the $m_{F}=0$ component, the k-space density is modulated obliquely as the local high densities are in that direction.
@@ III.2.2 Case-\@slowromancapii@: Cosine-Cosine magnetic field ${B}(\mathbf{r})=B_{0}(\cos x,\cos y)$
Figure 9: (a) shows the topological charge distribution, (b) shows the spin texture of the system and (c) shows the local amplification of the spin texture, marked by black square for the SO-coupled spin-1 BEC shown in Fig. 8. The color of the arrows represent the $S_{z}$ values. Units are dimensionless.
We present the associated topological charge distribution in Fig. 9 (a) and the spin texture is shown in Fig. 9 (b). For the diagonal vortices in the $m_{F}=0$ component, as the cores are symmetric, their respective skyrmion and anti-skyrmion structures are circular. However, for elongated vortices or anti-vortices, the corresponding topological charge distributions are also elliptical. We show in Fig. 9 (c), the local amplification of the spin texture for the region of the diagonal skyrmion and anti-skyrmion. The spin texture due to the cosine-cosine magnetic field has skyrmion structures along the oblique diagonal. Here, the anisotropy is induced in the direction corresponding to the phase difference in the applied magnetic field, as compared to the sine-sine case.
@@
It is paramount to note that in the absence of SOC, the ground state structures for the sine-sine case and the cosine-cosine case can be analogously explained by taking into account the fact that the topological excitations are observed for certain saddle-points in the magnetic fields. In this light, it becomes imperative to study the ground state structures for the sine-cosine and/or cosine-sine fields.
@@
Figure 10: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the sine-cosine magnetic field $B_{x}=B_{0}\sin x$ and $B_{y}=B_{0}\cos y$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
III.2.3 Case-\@slowromancapiii@: Sine-Cosine magnetic field $B(\mathbf{r})=B_{0}(\sin x,\cos y)$
In this case, we study the ground state structures for the spin-1 BEC with the external magnetic field chosen as $B_{x}=B_{0}\sin x$ and $B_{y}=B_{0}\cos y$. For this combination of magnetic fields, in the absence of SOC, the ground state structures exhibit density modulations due to the resultant interaction of the sine-cosine magnetic fields in the $x,y$ directions, respectively. We present the ground state density profiles of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components in Fig. 10
@@ III.2.3 Case-\@slowromancapiii@: Sine-Cosine magnetic field $B(\mathbf{r})=B_{0}(\sin x,\cos y)$
Figure 11: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-cosine magnetic field $B_{x}=B_{0}\sin x$ and $B_{y}=B_{0}\cos y$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
(a1-a3). In the absence of SOC, topological excitations in the form of vortices and anti-vortices are observed only in the $m_{F}=1$ and $m_{F}=-1$ components. These are arranged in pairs of alternating vorticity along the $y$-direction and as alternating vortex-anti-vortex triads in the $x$-direction. Due to different magnetic fields in different directions, the local high densities in the $m_{F}=0$ component (which arrise as an outcome of the spin-exchange interaction of $m_{F}=0$ component with the $m_{F}=\pm 1$ components), are anisotropic in nature. The vorticity profile (see Fig. 10 (b1-b3)) confirms the complementary response of the $m_{F}=1$ and $m_{F}=-1$ components to the external magnetic field with alternating vortices and anti-vortices. The momentum distribution is shown in Fig. 10 (c1-c3). For this case also, additional density peaks are observed in both the $k_{x}$ and $k_{y}$ directions in momentum space, alongside the central peak. However, as their magnitudes are not same in orthogonal directions, the density modulations have different wavelengths along different axes.
In the presence of spin-orbit coupling, the system undergoes symmetry breaking, manifesting in density modulations in opposing directions. Notably, these modulations exhibit distinct characteristics across different components of the system. The density as well as vortex distributions are highly anisotropic in nature due to the combined effect of SOC-induced anisotropy and different sinusoidal fileds in different directions. We present the component-wise density profiles in Fig. 11 (a1-a3). In the $m_{F}=1$ component, there is a local high density surrounded by vortices, which is compensated by vortices and vortex brights in the $m_{F}=-1$ component. The $m_{F}=0$ component has a unique configuration for which a local high density is observed, which is surrounded by anti-vortices along the $x$-direction and an adjacent vortex along the $y$-direction. The corresponding vorticity profiles are given in Fig. 11 (b1-b3). For this case also, the coreless vortices have deformed elliptical shape as a consequence of induced anisotropy in the system. The momentum distributions as shown in Fig. 11 (c1-c3) conform the effect of anisotropy in the density modulations in the spin-1 system. The k-space densities are elongated in opposite directions for the $m_{F}=1$ and $m_{F}=-1$ components, respectively, i.e., for the $m_{F}=1$ case, the central peak is slightly elongated along the $x-$axis while for the $m_{F}=-1$ component, the central density peak is slightly elongated in the $y-$direction. This is to respect the density modulations with different wavelengths in the anisotropic system. For the $m_{F}=0$ component, the momentum distribution is symmetric however not the same about the $k_{x},k_{y}$ axis.
The topological charge distribution demonstrated in Fig. 12 (a) reveals the presence of skyrmions and anti-skyrmions, which is confirmed by the spin texture of the system shown in Fig. 12 (b). Because of the combined effect of anisotropy and different magnetic fields along different directions, the spin texture has both skyrmions and anti-skyrmions. Owing to the vortex-anti-vortex triad, the skyrmion-anti-skyrmions are arranged in a triangular fashion. Fig. 12 (c) shows the local amplification of the spin texture showcasing the central vortex-anti-vortex triad.
@@
Figure 12: (a) shows the topological charge distribution, (b) shows the spin texture of the system and (c) shows the local amplification of the spin texture, marked by black square, for the SO-coupled spin-1 BEC shown in Fig. 11. The color of the arrows represent the $S_{z}$ values. Units are dimensionless.
III.2.4 Case-\@slowromancapiv@: Cosine-Sine magnetic field $B(\rm{r})=B_{0}(\cos x,\sin y)$
For the present case, we propose an alternate choice of cross-magnetic field in the $x$-$y$ plane, i.e., $B_{x}=B_{0}\cos x$ and $B_{y}=B_{0}\sin x$. Inspite of the striking similarities with the sine-cosine case (discussed in section III.2.3), there are subtle differences in the present case. It is trivial to understand that for this case also, the ground state structures are direction-dependent as a result of choosing different magnetic fields in different directions. Importantly, as the magnetic fields along the $x,y$-directions are swapped, their effect on the spin-1 ground state will therefore, be also swapped. In the absence of SOC, the density profiles, associated vorticities and momentum
Figure 13: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-cosine magnetic field $B_{x}=B_{0}\cos x$ and $B_{y}=B_{0}\sin y$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
Figure 14: (a) shows the topological charge distribution, (b) shows the spin texture of the system and (c) shows the local amplification of the spin texture, marked by black square, for the SO-coupled spin-1 BEC shown in Fig. 13. The color of the arrows represent the $S_{z}$ values. Units are dimensionless.
space densities of the three components are interchanged along the $x$-$y$ axes.
@@ III.2.4 Case-\@slowromancapiv@: Cosine-Sine magnetic field $B(\rm{r})=B_{0}(\cos x,\sin y)$
However, for the SO-coupled spin-1 BEC in the cosine-sine magnetic field, the ground state structures are not simply the transpose of that of the sine-cosine case. The ground state structures which are an outcome of the intricate interplay between the external magnetic field, SOC and the spinor order parameter is being presented in Fig. 13. The density profiles (see Fig. 13 (a1-a3)) vouch for the experienced anisotropy in the system due to SOC. The $m_{F}=1$ component has a unique configuration of vortices along with adjacent local high densities in the system. The $m_{F}=0$ component has a linear chain of vortices in the $y$-direction with an additional anti-vortex in the $x$-direction. The $m_{F}=-1$ component has a local high density surrounded by vortex structures on adjacent sides. The vorticity profile as shown in Fig. 13 (b1-b3) depicts the distribution of vortices or anti-vortices in the system. The coreless vortices have a slightly distorted shape due to the symmetry breaking. Fig. 13 (c1-c3) displays the momentum distributions. The presence of induced anisotropy in the system leads to observable elongations in the k-space densities along specific directions, owing to the varying periodicities of in-plane magnetic fields in the $x$ and $y$-directions. Moreover, notable density peaks are evident at certain non-zero $k_{x},k_{y}$.
@@
The topological charge distribution shown in Fig. 14 (a). The skyrmions and anti-skyrmions are observed in different axes. The spin texture is shown in Fig. 14 (b). Aside from the central triad formed with alternating vortex-anti-vortex configuration, a few other spin textures can be observed owing to the coreless vortices in the system. Fig. 14 (c) shows the local amplification of the spin texture showing the central triad in the system. As the sine-cosine and cosine-sine cases are not just mere transpose of each other in the presence of SOC, the skyrmion arrangement although in the form of triads are distinct from one another.
In this work, we have focussed our results for a weak SOC with strength $\gamma=0.5$, only. However, as we increase the $\gamma$ values, a very intricate arrangement of skyrmions and anti-skyrmions is formed due to the highly induced anisotropy in the system. As a consequence, the density modulations with short wavelengths are observed in the ground state configurations giving rise to complicated vortex-anti-vortex structures depending upon the choice of external in-plane magnetic field. We discuss the impact of variation of $\gamma$ on the spin texture in appendix A.
IV Conclusion